40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 21 18:18:03 2020, gautam, Update, CDS, Many model changes CDSoverview.png
    Reply  Tue Dec 22 22:11:37 2020, gautam, Update, CDS, AA filt repaired and reinstalled 
Message ID: 15744     Entry time: Tue Dec 22 22:11:37 2020     In reply to: 15743
Author: gautam 
Type: Update 
Category: CDS 
Subject: AA filt repaired and reinstalled 

Koji fixed the problematic channel - the issue was a bad solder joint on the input resistors to the THS4131. The board was re-installed. I also made a custom 2x4-pin LEMO-->DB9 cable, so we are now recording the PMC and FSS ERR/CTRL channel diagnostics again (spectra tomorrow). Note that Ch32 is recording some sort of DuoTone signal and so is not usable. This is due to a misconfiguration - ADC0 CH31 is the one which is supposed to be reserved for this timing signal, and not ADC1 as we currently have. When we swap the c1ioo hosts, we should fix this issue.

I also did most of the work to make the MEDM screens for the revised ASC topology, tried to mirror the site screens where possible. The overview screen remains to be done. I also loaded the anti-whitening filters (z:p 150:15) at the demodulated WFS input signal entry points. We don't have remote whitening switching capability at this time, so I'll test the switching manually at some point.

Quote:

The main issue is that in the AA chassis I built, Ch14 (with the first channel as Ch1) has the output saturated to 28V (differential). I'm not sure what kind of overvoltage protection the ADC has - we frequently have the inputs exceed the spec'd +/-20 V (e.g. when the whitening filters are engaged and the cavity is fringing), but pending further investigation, I am removing the SCSI connection from the rear of the AA chassis.

ELOG V3.1.3-